Pharaoh04:

Triangle XYZ was dilated by a scale factor of 2 to create triangle ACB and cos ∠X = 2 and 5 tenths over 5 and 59 hundredths. Triangles XYZ and ACB; angles Y and C both measure 90 degrees, angles A and X are congruent. Part A: Use complete sentences to explain the special relationship between the trigonometric ratios of triangles XYZ and ACB. You must show all work and calculations to receive full credit. (5 points) Part B: Explain how to find the measures of segments AC and AB. You must show all work and calculations to receive full credit. (5 points)

surjithayer:

 \[\angle X=\angle A (given)\] \[\angle Y= \angle C (each=90 ^\circ)\] \[\Delta ACB \cong \Delta XYZ\] so sides are proportional.

surjithayer:

\[\frac{ XY }{ AC }=\frac{ YZ }{ CB}=\frac{ XZ }{AB }\] OR \[\frac{ XY }{ XZ }=\frac{ AC }{ AB }\] \[\cos X=\frac{ XY }{ XZ }=\frac{ AC }{ AB}=\cos C\] cos X=cos C similarly other trigonometric ratios are also equal.